OpenStudy (hy123):

4x^2-6=5x solve for x

OpenStudy (anonymous):

4x^2-5x-6=0; D=25+4*4*6=121; x1= (5+11)/2*4=2; x2=(5-11)/2*4=-0.75

OpenStudy (hy123):

so x= 2,-0.75?

OpenStudy (anonymous):

yes

OpenStudy (hy123):

ty